OpenStudy (anonymous):

Solve for x: –4(3x – 2) = 6x + 2

OpenStudy (anonymous):

are you really having trouble with that because its kinda easy

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(-4)(3x) - (-4)(2) = 6x + 2 -12x - (-8) = 6x + 2 -12x + 8 = 6x + 2 bring all the x terms to one side of the equation 8 - 2 = 6x + 12x 6 = 18x 6/18 =x 1/3 =x

OpenStudy (petiteme):

–4(3x – 2) = 6x + 2 distribute -4 -12x + 8 = 6x + 2 transpose -12x to the other side as well as 2 8-2 = 6x +12x do the mathematics :) 18x=6 divide both sides by 18 :) x=6/18 or 1/3

OpenStudy (anonymous):

just give me the damn answer

OpenStudy (anonymous):

try not to fail in your math test though. you're welcome